# AttackSeqBench: Benchmarking Large Language Models in Analyzing Attack Sequences within Cyber Threat Intelligence

### v2

@@ 5 Conclusion
The breakthrough of LLMs has shown promising potential across the cybersecurity domain, particularly in CTI understanding. Despite this, the applicability of LLMs in analyzing adversarial sequences remains largely unexplored. In this work, we propose AttackSeqBench, a benchmark tailored for assessing LLMs’ ability in understanding how adversaries operate through inferring TTPs based on attack sequences from real-world CTI reports. To cater to the evolving threat landscape, we design an automated Q&A construction pipeline that enables the Extensibility of our benchmark to new CTI reports. We further conduct extensive experiments across three settings with varying context availability, evaluating diverse LLMs, LRMs, and post-training strategies to verify its Reasoning Scalability and Domain-Specific Epistemic Expandability and thoroughly analyze their ability boundaries in attack sequence analysis. Our work opens up a new direction towards LLM-driven CRU, enabling effective threat intelligence mining through automation.
@@
Reproducibility Statement
To promote reproducibility, we release an anonymous repository that contains all resources necessary to replicate our study, including the original CTI reports, the dataset construction pipeline, the complete datasets for the three tasks (i.e., AttackSeq-Tactic, AttackSeq-Technique, and AttackSeq-Procedure), and the code for running the three benchmark settings (i.e., Zero-Shot setting, Context setting, and RAG-empowered setting). In addition, we also provide the comprehensive implementation details about the four post-training strategies in Appendix. Together, these resources ensure that our AttackSeqBench can be reliably reproduced and the reported results independently verified.
Appendices
@@ Appendix A Dataset
A.1 Dataset Distribution
Based on Figure 7, we observe that the top three most frequent tactics (i.e., Command and Control, Defense Evasion and Execution) occur in the middle of attack sequences, while the bottom two tactics (i.e., Exfiltration and Reconnaissance) occurs at the start and the end of the attack sequence. Similarly, the most frequent ATT&CK technique is T1071-Application Layer Protocol  [^3] https://attack.mitre.org/techniques/T1071/, which is associated with the most common operations of APTs, Command and Control.
Figure 7: Visualization of the distribution of top-10 tactics and techniques in AttackSeqBench.
A.2 Dataset Evaluation Criteria
@@
To quantitatively evaluate its quality, we first design the 5-point Likert scale for each aspect (refer to Table LABEL:tab.annotation_instructions), where each score corresponds to a different level of the given aspect. Then we instruct three cybersecurity experts and LLM to provide the score of each aspect to achieve the human evaluation and the automatic evaluation, respectively. The detailed results are shown in Table 1. While the automatic evaluation results are lower than human evaluation, the human evaluation shows that most Q&A pairs in the dataset satisfy the requirements of all aspects. This suggests that automatic evaluation is still limited in knowledge-intensive domains such as in cybersecurity. Note that for the AttackSeq-Procedure-No, we evaluate questions only on four aspects—Answerability, Clarity, Consistency, and Answer Consistency—since it is derived from AttackSeq-Procedure-Yes through negation of temporal prepositions and replacement of procedures.
A.3 Benchmark Tasks
@@
Table 5: Annotation instructions for each evaluation aspect.
@@
Consistency Score 2: The question loosely aligns with the TTP description but has inconsistencies or inaccuracies.
@@
AttackSeq-Tactic. This task evaluates the LLMs’ capability to infer a tactic $t_{k}\in T$. Given a question $Q$ that corresponds to tactic $t_{k}$ and four shuffled candidate tactics $C_{Tac}=\{c_{r}:r\in[1,4]\}$, the LLM will be instructed to select the correct tactic $c_{l}\in C_{Tac}$.
AttackSeq-Technique. This task assesses the LLMs’ capability to infer a technique $e_{j,k}\in E(t_{k})$. Given a question $Q$ that corresponds to $e_{j,k}$ and four shuffled candidate techniques $C_{Tec}=\{c_{r}:r\in[1,4]\}$, the LLM will be instructed to select the correct technique $c_{l}\in C_{Tec}$.
AttackSeq-Procedure. This task challenges the LLMs’ capability to determine the likelihood of procedures $p_{m,j,k}\in P(e_{j,k})$ in an attack sequence. Given a question $Q$ and two candidate choices $C_{Pro}=\{yes,no\}$, the LLM will be instructed to determine if the procedure $p_{m,j,k}$ is likely to occur in the given attack sequence $S$.
We further divide AttackSeq-Procedure into two sub-tasks, namely AttackSeq-Procedure-Yes and AttackSeq-Procedure-No, based on the ground truth of the boolean question. This explores the LLMs’ ability in determining misleading procedures that are unlikely to occur in an attack sequence.
A.4 Baselines
To demonstrate the effectiveness and robustness of our proposed AttackSeqBench, we evaluate seven large language models, five large reasoning models and four post-training strategies across three tasks involving different levels of data and three benchmark settings with varying context completeness. We leverage vLLM (Kwon et al., 2023) to run all the open-source LLMs locally with two Nvidia H100 GPUs. For the colsed-source LLMs, we utilize OpenAI’s Batch API  [^4] https://platform.openai.com/docs/guides/batch to conduct inference in batches. In our experiments, we set the following sampling parameters while keeping the default value for the remaining parameters: temperature to 0, maximum output tokens to 2048, and top_p to 1. Below is the details of the utilized LLMs, LRMs and post-training strategies in our experiments:
@@ A.4 Baselines
  • LLaMa-3.1-8B (Grattafiori et al., 2024) is an instruction-tuned LLM from Meta, balancing performance and efficiency for textual understanding tasks.
  • ChatGLM-4-9B (GLM et al., 2024) is pretrained on ten trillions of tokens and further achieve the high-quality alignment through supervised fine-tuning and human feedback learning.
  • Qwen2.5-3B, Qwen2.5-14B, Qwen2.5-32B and Qwen2.5-72B (QwenTeam, 2024) represents Qwen-2.5 series LLMs with different parameter scales, demonstrating strong instruction-following and long-text generation capabilities.
  • LLaMa-3.3-70B (Grattafiori et al., 2024) is an auto-regressive language model which is instruction-tuned in 70B with SFT and reinforcement learning with human feedback (RLHF).
@@
  • DeepSeek-R1-Distill-Llama-8B (R1 (Llama-8B)), DeepSeek-R1-Distill-Qwen-14B (R1 (Qwen-14B)) and DeepSeek-R1-Distill-Qwen-32B (R1 (Qwen-32B)) (DeepSeek-AI, 2025) are fine-tuned from the Llama-3.1-8B, Qwen2.5-14B and Qwen2.5-32B with 800k samples curated with DeepSeek-R1, aiming to equip smaller models with reasoning capabilities like DeepSeek-R1.
@@
A.5 Post-training Corpus Construction
Considering the post-training strategies, we construct two diverse datasets for SFT and RD, RLIF and RLVR respectively. For the former, we utilize a subset of the Primus-Instruct dataset (Yu et al., 2025). Primus-Instruct is a cybersecurity corpus collected for instruction-tuning, containing diverse task types such as alert explanation, suspicious command analysis, security event query generation, retrieved security document QA, Terraform security mis-configuration repair, and general multi-turn instruction following. To mitigate the inherent bias from linguistic inconsistencies, we filter out non-English samples via the FastText language identification library (Joulin et al., 2016) and manually verify the results, yielding a subset of 710 samples for SFT.
@@ A.6 Implementation Details
Retrieval Augmented Generation (RAG): We first crawl the description and the example procedures of each technique in the Enterprise ATT&CK Matrix v17  [^5] https://attack.mitre.org/versions/v17/matrices/enterprise/. Then, we split the textual data into text chunks and embed the chunks in a vector store (i.e., Chroma DB  [^6] https://www.trychroma.com/), where each chunk’s metadata contains the associated ATT&CK tactic and technique. We utilize a hybrid retriever with a re-ranker, by combining combine Okapi BM25 (Robertson & Zaragoza, 2009) and a dense retriever based on the more advanced text-embedding-3-large from OpenAI (OpenAI, 2024b). We set the chunk size to 512 and retrieved chunks to 3, and utilize LlamaIndex (Liu, 2022) to implement the retriever. Additionally, we also implement BGE-EN-ICL (Li et al., 2024) and ATT&CK-BERT (Abdeen et al., 2023a) within RAG to evaluate their effectiveness in attack sequence analysis.
@@
Reasoning Distillation (RD) refers to fine-tune LLM on a reasoning dataset distilled from the advanced LRMs (i.e., DeepSeek-R1  and GPT-o1-preview ), which enables the smaller LLM to inherit the reasoning behaviors of the above LRMs. For RD, we fine-tune our backbone LLM on the latter dataset within Appendix A.5 with the same parameter settings in SFT.
Reinforcement Learning with Verifiable Rewards (RLVR) extends reinforcement learning by incorporating verifiable signals as rewards, such as correctness checks or logical consistency that can be programmatically validated. Specifically, we implement RLVR with Group Relative Policy Optimization (GRPO) (Shao et al., 2024) on the Volcano Engine Reinforcement Learning (verl) framework, where group-normalized rewards reduce variance and stabilize training. We conduct RLVR with the learning rate of $3\times 10^{-6}$ using Fully Sharded Data Parallel (FSDP).
@@
Figure 8: Comparison between existing CTI-related benchmarks and our AttackSeqBench based on a real-world CTI report (Koehl & Hannon, 2020). Our benchmark emphasizes on the sequences of adversary behaviors described in CTI reports.
A.7 Related Benchmarks Comparison
@@
A.8 The Use of Large Language Models
We declare that LLMs were employed to assist with the refinement of this manuscript, specifically for grammar checking and language polishing. Additionally, LLMs were used in a limited capacity for minor debugging and syntactic correction of code snippets. Beyond these auxiliary roles, given that the primary purpose of this work is to explore the capability of diverse types of LLMs in understanding adversarial attack sequences, we also utilize a range of open-source and closed-source LLMs and LRMs during dataset construction, dataset refinement, and performance evaluation. All such uses are documented in the main paper and the appendix and were carefully controlled to ensure transparency and reproducibility.
A.9 Limitations and Future works
Limitations: While our work serves as a pioneering study into the LLMs’ reasoning capabilities in attack sequence analysis, several limitations should be acknowledged. Firstly, our study focuses on correctness of models’ responses through Multi-Choice Questions and Yes-No Questions, which may not fully capture the reasoning abilities of LLMs necessary for comprehensive evaluation. Secondly, although we have conducted extensive experiments with seven LLMs, five LRMs, and four post-training strategies across three benchmark tasks (AttackSeq-Tactic, AttackSeq-Technique, and AttackSeq-Procedure) and three benchmark settings (Zero-Shot setting, Context setting, and RAG-empowered setting), fully demonstrating the Reasoning Scalability and Domain-Specific Epistemic Expandability of our AttackSeqBench, the implementations of RAG and post-training strategies remains relatively basic and leave room for future refinement. Thirdly, our AttackSeqBench currently leverages 408 rigorously filtered CTI reports to extract attack sequences and generate Q&A pairs. Although this number substantially exceeds prior CTI-related studies (i.e., 12 in AttacKG+ (Zhang et al., 2025c), 12 in MM-AttacKG (Zhang et al., 2025d), and at most 71 in Attack Flow  [^7] https://center-for-threat-informed-defense.github.io/attack-flow/), the proposed dataset construction pipeline is flexible and can be readily extended to unseen CTI reports. This not only demonstrates the Extensibility of our AttackSeqBench but also highlights an important direction for continuously refining this benchmark in future work. Nevertheless, while it is important to be aware of these limitations, our AttackSeqBench serves are a valuable benchmark to systematically explore LLMs’ reasoning abilities across the tactical, technical and, procedural dimensions of adversarial behaviors.
Future works: Building on the limitations, our future research on AttackSeqBench will proceed along three directions. Regarding evaluation, we plan to expand our evaluation methods from the simple Multiple-Choice Question tasks and Yes-No Question tasks to the more complex reasoning and completion tasks, thereby providing a more comprehensive assessment of model capabilities in CTI report understanding. In terms of methodology, we will build on AttackSeqBench to explore more fine-grained RAG approaches and advanced post-training strategies that account for the knowledge-extensive and high-stakes nature of CTI reports understanding, aiming to fully leverage model potential in complex cyber-attack scenarios. At the data level, we will continue to expand and dynamically update the CTI corpus to ensure our AttackSeqBench remains evolvable over time, thereby supporting the steady advancement of domain-specific foundation models for cybersecurity.
@@ Appendix B Experiments
B.1 TTP Temporal Position Analysis
To better understand the LLMs’ capabilities in attack sequence analysis, we conduct fine-grained analysis based on each stage within the attack sequences of MITRE ATT&CK®. We illustrate the performance of two LLMs and two LRMs across all benchmark tasks in the Regular setting in Figure 9 and show the corresponding values, the mean and standard deviation (SD) of these LLMs and LRMs on each tactic and benchmark task in Table 6. We identify four overachieving attack phases to categorize the ATT&CK tactics in attack sequences: (1) Initial Intrusion Phase; (2) Exploitation Phase; (3) Stealth Expansion Phase; (4) Objective Orchestration Phase. It is worth noting that our categorization follows Tactics, as each Technique and Procedure in MITRE ATT&CK® is uniquely mapped to a specific Tactic within a given attack sequence.
Figure 9: The performance comparison between two LLMs and two LRMs on each tactic in each benchmark task.
Table 6: Results of performance of two LLMs and two LRMs on all the 14 tactics, and the corresponding statistics of mean and standard deviation.
Tactic

### v3

@@ 5. Conclusion
The breakthrough of LLMs has shown promising potential across the cybersecurity domain, particularly in CTI analysis. Despite this, the applicability of LLMs in understanding attack sequences remains largely unexplored. In this work, we propose AttackSeqBench, a benchmark tailored for assessing LLMs’ ability in understanding how adversaries operate through inferring TTPs based on attack sequences from real-world CTI reports. To cater to the evolving threat landscape, we design an automated Q&A construction pipeline that enables the Extensibility of our benchmark to new CTI reports. We further conduct extensive experiments across three settings with varying context availability, evaluating diverse LLMs, LRMs, and post-training strategies to verify its Reasoning Scalability and Domain-Specific Epistemic Expandability and thoroughly analyze their ability boundaries in attack sequence analysis. Our work opens up a new direction towards LLM-driven CRU, enabling effective threat intelligence mining through automation.
6. Limitations and Future works
Limitations: While our work serves as a pioneering study in benchmarking LLMs’ capability in attack sequence understanding, several limitations should be acknowledged. Firstly, our study focuses on the correctness of LLMs’ responses through Multi-Choice Questions and Yes-No Questions, which may not fully capture the reasoning abilities of LLMs. Secondly, although we have conducted extensive experiments with seven LLMs, five LRMs, and four post-training strategies across three benchmark tasks (AttackSeqBench-Tactic, AttackSeqBench-Technique, and AttackSeqBench-Procedure) and three benchmark settings (Zero-Shot setting, Context setting, and RAG-empowered setting), fully demonstrating the Reasoning Scalability and Domain-Specific Epistemic Expandability of our AttackSeqBench, the implementations of RAG and post-training strategies remains relatively basic and leave room for future refinement. Thirdly, our AttackSeqBench currently leverages 408 rigorously filtered CTI reports to extract attack sequences and generate Q&A pairs. Although this number substantially exceeds prior CTI-related studies (i.e., 12 in AttacKG+ (Zhang et al., 2025c)  [^4] Notably, AttacKG+ just evaluates on only 12 CTI reports, the “500 CTI reports” are merely used to show extracted TTP distributions rather than method effectiveness., 12 in $M$-AttacKG (Zhang et al., 2026), and at most 71 in Attack Flow  [^5] https://center-for-threat-informed-defense.github.io/attack-flow/), the proposed Q&A dataset construction pipeline (source codes are publicly-available in our GitHub repo) is flexible and can be readily extended to unseen CTI reports. This not only demonstrates the Extensibility of our AttackSeqBench but also highlights an important direction for continuously refining this benchmark in future work. Nevertheless, while it is important to be aware of these limitations, our AttackSeqBench serves as a valuable benchmark to systematically explore LLMs’ reasoning abilities across the tactical, technical and, procedural dimensions of adversarial behaviors.
Future works: Building on the limitations, our future research on AttackSeqBench will proceed along three directions. Regarding evaluation, we plan to expand our evaluation methods from the simple Multiple-Choice Question tasks and Yes-No Question tasks to the more complex reasoning and completion tasks, thereby providing a more comprehensive assessment of LLM capability in attack sequence understanding. In terms of methodology, we will build on AttackSeqBench to explore more fine-grained RAG approaches and advanced post-training strategies that account for the knowledge-extensive and high-stakes nature of CTI reports, aiming to fully leverage model potential in complex cyber-attack scenarios. At the data level, we will continue to expand and dynamically update the CTI corpus to ensure our AttackSeqBench remains evolvable over time, thereby supporting the steady advancement of domain-specific foundation models for cybersecurity.
@@ Appendix A Dataset
A.1. Dataset Distribution
Based on Figure 7, we observe that the top-3 most frequent tactics (i.e., Command and Control, Defense Evasion and Execution) occur in the middle of attack sequences, while the bottom two tactics (i.e., Exfiltration and Reconnaissance) occurs at the start and the end of the attack sequence. Similarly, the most frequent ATT&CK technique is T1071-Application Layer Protocol  [^6] https://attack.mitre.org/techniques/T1071/, which is associated with the most common operations of APTs, Command and Control.
Figure 7. Visualization of the distribution of top-10 tactics and techniques in AttackSeqBench.
A.2. Dataset Evaluation Criteria
@@
To quantitatively evaluate its quality, we first design the 5-point Likert scale for each aspect (refer to Table 5), where each score corresponds to a different level of the given aspect. Then we instruct three cybersecurity experts and LLM to provide the score of each aspect to achieve the human evaluation and the automatic evaluation, respectively. The detailed results are shown in Table 1. While the automatic evaluation results are lower than human evaluation, the human evaluation shows that most Q&A pairs in the dataset satisfy the requirements of all aspects. This suggests that automatic evaluation is still limited in knowledge-intensive domains such as in cybersecurity. Note that for the AttackSeqBench-Procedure-No, we evaluate questions only on four aspects—Answerability, Clarity, Consistency, and Answer Consistency—since it is derived from AttackSeqBench-Procedure-Yes through negation of temporal prepositions and replacement of procedures.
A.3. Benchmark Tasks
@@
Table 5. Annotation instructions for each evaluation aspect.
@@
Score 2: The question loosely aligns with the TTP description but has inconsistencies or inaccuracies.
@@
AttackSeqBench-Tactic. This task evaluates the LLMs’ capability to infer a tactic $t_{k}\in T$. Given a question $Q$ that corresponds to tactic $t_{k}$ and four shuffled candidate tactics $C_{Tac}=\{c_{r}:r\in[1,4]\}$, the LLM will be instructed to select the correct tactic $c_{l}\in C_{Tac}$.
AttackSeqBench-Technique. This task assesses the LLMs’ capability to infer a technique $e_{j,k}\in E(t_{k})$. Given a question $Q$ that corresponds to $e_{j,k}$ and four shuffled candidate techniques $C_{Tec}=\{c_{r}:r\in[1,4]\}$, the LLM will be instructed to select the correct technique $c_{l}\in C_{Tec}$.
AttackSeqBench-Procedure. This task challenges the LLMs’ capability to determine the likelihood of procedures $p_{m,j,k}\in P(e_{j,k})$ in an attack sequence. Given a question $Q$ and two candidate choices $C_{Pro}=\{yes,no\}$, the LLM will be instructed to determine if the procedure $p_{m,j,k}$ is likely to occur in the given attack sequence $S$.
Given the AttackSeqBench-Procedure, we further divide it into two sub-tasks based on the ground truth of the boolean question, namely AttackSeqBench-Procedure-Yes and AttackSeqBench-Procedure-No. This explores the LLMs’ ability in determining misleading procedures that are unlikely to occur in an attack sequence.
A.4. Baselines
To demonstrate the effectiveness and robustness of our proposed AttackSeqBench, we evaluate seven large language models, five large reasoning models and four post-training strategies across three tasks involving different levels of data and three benchmark settings with varying context completeness. We leverage vLLM (Kwon et al., 2023) to run all the open-source LLMs locally with two Nvidia H100 GPUs. For the colsed-source LLMs, we utilize OpenAI’s Batch API  [^7] https://platform.openai.com/docs/guides/batch to conduct inference in batches. In our experiments, we set the following sampling parameters while keeping the default value for the remaining parameters: temperature to 0, maximum output tokens to 2048, and top_p to 1. Below are the details of the utilized LLMs, LRMs and post-training strategies in our experiments:
@@ A.4. Baselines
  • LLaMa3.1-8B (Grattafiori et al., 2024) is an instruction-tuned LLM from Meta, balancing performance and efficiency for textual understanding tasks.
  • ChatGLM4-9B (GLM et al., 2024) is pretrained on ten trillions of tokens and further achieve the high-quality alignment through supervised fine-tuning and human feedback learning.
  • Qwen2.5-3B, Qwen2.5-14B, Qwen2.5-32B and Qwen2.5-72B (QwenTeam, 2024) represents Qwen2.5 series LLMs with different parameter scales, demonstrating strong instruction-following and long-text generation capabilities.
  • LLaMa3.3-70B (Grattafiori et al., 2024) is an auto-regressive language model which is instruction-tuned in 70B with SFT and reinforcement learning with human feedback (RLHF).
@@
  • DeepSeek-R1-Distill-Llama-8B (R1 (LLaMa-8B)), DeepSeek-R1-Distill-Qwen-14B (R1 (Qwen-14B)) and DeepSeek-R1-Distill-Qwen-32B (R1 (Qwen-32B)) (DeepSeek-AI, 2025) are fine-tuned from the LLaMa3.1-8B, Qwen2.5-14B and Qwen2.5-32B with 800k samples curated with DeepSeek-R1, aiming to equip smaller models with reasoning capabilities like DeepSeek-R1.
@@
A.5. Post-training Corpus Construction
Considering the post-training strategies, we construct two diverse datasets for instruction-tuning (i.e., SFT) and reinforcement learning (i.e., RLVR, RD and RLIF) respectively. For the former, we utilize a subset of the Primus-Instruct dataset (Yu et al., 2025). Primus-Instruct is a cybersecurity corpus collected for instruction-tuning, containing diverse task types such as alert explanation, suspicious command analysis, security event query generation, retrieved security document Q&A, Terraform security mis-configuration repair, and general multi-turn instruction following. To mitigate the inherent bias from linguistic inconsistencies, we filter out non-English samples via the FastText language identification library (Joulin et al., 2016) and manually verify the results, yielding a subset of 710 samples for SFT.
@@ A.6. Implementation Details
Retrieval Augmented Generation (RAG): We first crawl the description and the example procedures of each technique in the Enterprise ATT&CK Matrix v17  [^8] https://attack.mitre.org/versions/v17/matrices/enterprise/. Although alternative threat modeling approaches such as the Diamond Model and the Lockheed Martin Kill Chain exist, they are fundamentally abstractions that constrain cyber attacks, differing primarily in their modeling emphases. Here, we adopt the ATT&CK framework, as it is continuously maintained and underpins the majority of existing TTP extraction efforts. Then, we split the textual data into several chunks and embed them in a vector store (i.e., Chroma DB  [^9] https://www.trychroma.com/), where each chunk’s metadata contains the associated ATT&CK tactic and technique. We utilize a hybrid retriever with a re-ranker, by combining BM25 (Robertson and Zaragoza, 2009) as the sparse retriever and a dense retriever based on the more advanced text-embedding-3-large from OpenAI (OpenAI, 2024b). We set the chunk size to 512 and retrieved chunks to 3, and utilize LlamaIndex (Liu, 2022) to implement the retriever. Additionally, we also implement BGE-EN-ICL (Li et al., 2024) and ATT&CK-BERT (Abdeen et al., 2023b) within RAG to evaluate their effectiveness in attack sequence analysis.
@@
Reasoning Distillation (RD) refers to fine-tune LLM on a reasoning dataset distilled from the advanced LRMs (i.e., DeepSeek-R1 (DeepSeek-AI, 2025) and GPT-o1-preview (OpenAI, 2024a)), which enables the smaller LLM to inherit the reasoning behaviors of the above LRMs. For RD, we fine-tune our backbone LLM on the latter dataset within Appendix A.5 with the same parameter settings in SFT.
@@
Reinforcement Learning with Verifiable Rewards (RLVR) extends reinforcement learning by incorporating verifiable signals as rewards, such as correctness checks or logical consistency that can be programmatically validated. Specifically, we implement RLVR with Group Relative Policy Optimization (GRPO) (Shao et al., 2024) on the Volcano Engine Reinforcement Learning (verl) framework, where the utilized reward combines an accuracy reward and a format reward for each query, and group-normalized rewards reduce variance and stabilize training. We conduct RLVR with the learning rate of $3\times 10^{-6}$ using Fully Sharded Data Parallel (FSDP).
Figure 8. Comparison between existing CTI-related benchmarks and our AttackSeqBench based on a real-world CTI report (Koehl and Hannon, 2020). Our benchmark emphasizes on the sequences of adversary behaviors described in CTI reports.
A.7. Related Benchmarks Comparison
@@ Appendix B Experiments
B.1. TTP Temporal Position Analysis
Figure 9. The performance comparison between two LLMs and two LRMs on each tactic in each benchmark task.
Table 6. Results of performance of two LLMs and two LRMs on all the 14 tactics, and the corresponding statistics of mean and standard deviation.
Tactics